OpenStudy (anonymous):

My friend has $104 to spend on a fence for her rectangular garden. She wants to use cedar fencing which costs $9/yard on one side, and cheaper metal fencing which costs $4/yard for the other three sides. What are the dimensions of the garden with the largest area she can enclose?

OpenStudy (anonymous):

thats kind of a tricky one

OpenStudy (anonymous):

i agree. do you have any idea how to start it?

OpenStudy (anonymous):

hmmm I would just start drawing pictures and start making some calculations

OpenStudy (anonymous):

i have 9y+4y+8x=104 i am just not sure like what to plug it into to turn it into a quadratic

OpenStudy (anonymous):

whats 8x?

OpenStudy (anonymous):

the two metal sides. since it is a rectangle i thought that. y=length, while x=width

OpenStudy (anonymous):

sorry, should be 13y + 8x = 104

OpenStudy (anonymous):

so yes, you were correct.

OpenStudy (anonymous):

How would i find the max peremiter or area. plugged it into xy=a but that is at the origin.

OpenStudy (anonymous):

0=-1.625y^2+13y

OpenStudy (mertsj):

The largest area will be a square.

OpenStudy (anonymous):

the problem asks for a rectangle however, will the largest area always be in a square?

OpenStudy (mertsj):

yes

OpenStudy (mertsj):

a square is a rectangle.

OpenStudy (anonymous):

so just to get this straight, when a problem asks for max area, alaways find a square? and how i use the equation i solved for y to find the area or perimeter?

OpenStudy (mertsj):

For a given perimeter, the maximum area is that of a square. For your problem: 9x+12x=104 where x is a side of the square.

OpenStudy (anonymous):

i see, because y=x because it is a square?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

is 104/21 correct for a side?

OpenStudy (mertsj):

That's what I got.

OpenStudy (anonymous):

9x=44.57 12x=59.43 but it says this is wrong

OpenStudy (anonymous):

actually nevermind that is the price not length

OpenStudy (anonymous):

it actually still says that is incorrect. it says to round to 2 decimal places and iput 4.95

OpenStudy (mertsj):

21x=104 x=4.95 Area = x^2=24.53

OpenStudy (mertsj):

4.95 yd by 4.95 yd are the dimensions

OpenStudy (anonymous):

i know, i just emailed my teacher because all of the math add up, thank you for your help

OpenStudy (mertsj):

yw